### v1

# CroMe: Multimodal Fake News Detection using Cross-Modal Tri-Transformer and Metric Learning

Abstract
Multimodal Fake News Detection has received increasing attention recently. Existing methods rely on independently encoded unimodal data and overlook the advantages of capturing intra-modality relationships and integrating inter-modal similarities using advanced techniques. To address these issues, Cross-Modal Tri-Transformer and Metric Learning for Multimodal Fake News Detection (CroMe) is proposed. CroMe utilizes Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models (BLIP2) as encoders to capture detailed text, image and combined image-text representations. The metric learning module employs a proxy anchor method to capture intra-modality relationships while the feature fusion module uses a Cross-Modal and Tri-Transformer for effective integration. The final fake news detector processes the fused features through a classifier to predict the authenticity of the content. Experiments on datasets show that CroMe excels in multimodal fake news detection.
1 Introduction
Advances in information technology and the growing use of social media have made these platforms central to information exchange Mitra et al. (2017). However, this has also led to a rise in fake news, which harms public opinion, disrupts political stability Shu et al. (2017), and affects social and economic activities Xia et al. (2023). Detecting false information is essential to prevent its spread. The shift to digital platforms has changed how people consume news, increasing exposure to misinformation. Traditional methods, such as identifying logical flaws or obvious signs like spelling errors and image alterations, were effective for text or image content. However, the rise of multimedia formats like images and videos has accelerated the spread of fake news Jin et al. (2017), making multimedia-focused detection methods necessary.
Figure 1 shows four examples from the Weibo and Weibo-21 datasets. (a) A group of Quora users listed historically ingenious but unreasonable weapons which are semantically contradictory. (b) In Marco Island, dolphins rescued a dog from a canal and firefighters reunited it with its owner which are semantically conflicting. (c) Yichang police seized 7 macaques and arrested 2 suspects which is unrelated to the tiger in the image. (d) The text mentions abandoned Douglas DC 3 of the Aurora’s light but the image shows only the Aurora’s light. The main challenge in multimodal fake news detection is effectively identifying incongruent semantic features in both intra-modality and inter-modality contexts.
(a) A group of Quora users listed historically ingenious but unreasonable weapons.
(b) In Marco Island, dolphins rescued a dog and firefighters reunited it with its owner.
(c) Yichang police seized 7 macaques and arrested 2 suspects.
(d) Photographer captured an abandoned Douglas DC 3 under Aurora’s light.
Figure 1: Some fake examples from the Weibo and Weibo-21 datasets include: (a) The image shows intra-modality inconsistency. (b) Both the image and text contain intra-modality inconsistencies. (c) The image and text are unrelated. (d) The image does not match the text.
Fake news detection has progressed from text-based methods to advanced deep learning techniques. Early work focused on text analysis Castillo et al. (2011), while later studies utilized deep neural networks for linguistic and temporal patterns Ma et al. (2016), and attention mechanisms with RNNs Chen et al. (2018). Jin et al. Jin et al. (2016) introduced multimodal approaches by combining image, text, and social context features. Choi et al. Choi and Kim (2024) developed TT-BLIP, integrating text, image, and multimodal features using BLIP Li et al. (2022) encoders and a MultiModal Tri-Transformer for feature fusion. However, it primarily focuses on modality combination without addressing inter-modal and intra-modal relationships.
This paper introduces a Multimodal Fake News Detection approach using the Cross-Modal Tri-Transformer and Metric Learning (CroMe) network, a motivated framework inspired by the TT-BLIP model. Text is processed using pre-trained BERT Devlin et al. (2019), and images are analyzed with Masked Autoencoders (MAE) He et al. (2022). The BLIP2 Li et al. (2023) model captures semantic features across modalities. Additionally, CroMe introduces Cross-Modal fusion Chen et al. (2022) to compute and integrate cross-modal similarities, addressing the limitations of existing methods in capturing complex inter-modal interactions. For intra-modal relationships, a type of metric learning Schroff et al. (2015); Movshovitz-Attias et al. (2017) called the Proxy Anchor Loss method Kim et al. (2020) is used. This method captures fine-grained similarities while maintaining the efficiency of proxy-based methods through dynamic gradients. Thus, CroMe ensures that features within the same modality are more closely aligned while capturing and integrating cross-modal correlations, enhancing the overall feature representation for more accurate fake news detection. Experiments performed using the Weibo and Weibo-21 datasets demonstrate that CroMe outperforms state-of-the-art models, highlighting its effectiveness in Multimodal Fake News Detection. For the Politifact dataset, CroMe achieved results comparable to state-of-the-art models due to its smaller size.
This paper offers four key contributions:
  • The proposed multimodal fake news detection model, called CroMe, utilizes the advanced BLIP2 model for feature extraction to capture detailed text, image, and combined image-text features.
  • The model uses a proxy anchor method in the metric learning module to capture intra-modality relationships ensuring effective representation learning within the same modality.
  • The Cross-Modal Tri-Transformer Fusion, a new fusion technique, is introduced to calculate and merge similarities between text, image and image-text embeddings enhancing the model’s ability to handle interactions among different modalities.
  • CroMe is evaluated using the Weibo and Weibo-21 datasets, demonstrating that it outperforms state-of-the-art models and highlights its effectiveness in multimodal fake news detection. For the Politifact dataset, CroMe achieved comparable results to state-of-the-art models, likely due to the dataset’s smaller size.
2 Related work
Figure 2: Overview of the CroMe architecture. Masked Autoencoders (MAE), BERT, and BLIP2 encode multimodal features. Metric Learning extracts intra-modal relationships by representing class data points with proxies, where arrow thickness indicates the gradient scale. The CMTTF and Fake News Detector modules use cross-modal fusion and fake news detection.
Several methods for multimodal fake news detection focus on extracting features from text and images. EANN Wang et al. (2018) uses an event discriminator, and MCAN Wu et al. (2021) integrates textual and visual features through co-attention layers. MVAE Khattar et al. (2019) employs a multimodal variational autoencoder to reconstruct text and image features. Spotfake Singhal et al. (2019) applies BERT Devlin et al. (2019) for text and VGG19 Simonyan and Zisserman (2014) for images, with Spotfake+ Singhal et al. (2020) extending this to full articles. SAFE Zhou et al. (2020) detects fake news by analyzing text and visual similarities, while CAFE Chen et al. (2022) measures cross-modal ambiguity using KL divergence Kullback and Leibler (1951). LIIMR Singhal et al. (2022) emphasizes primary modalities and minimizes less significant ones. DistilBert Allein et al. (2021) detects disinformation by analyzing user preferences and sharing behaviors. BDANN Zhang et al. (2020) combines multimodal features to reduce event-specific biases. FND-CLIP Zhou et al. (2023) uses the CLIP model to measure image-text correlations and aggregate features through modality-wise attention. TT-BLIP Choi and Kim (2024) employs BLIP Li et al. (2022) encoders for text, image, and multimodal data, with the Multimodal Tri-Transformer fusing features using multi-head attention mechanisms.
Unlike previous models, the proposed model integrates uni-modal features and cross-modal correlations. The metric learning module aligns features within the same modality, while cross-modal similarities capture inter-modality interactions. By focusing on both intra-modality alignment and inter-modality integration, this approach improves fake news detection in social media.
3 Methodology
(a) Proxy anchor loss
(b) Triplet loss
Figure 3: Comparison of two different metric learning methods; (a) Proxy as an anchor, and (b) Data point as an anchor. The thickness of arrows in proxy anchor loss indicates the gradient scale determined by the scaling factor $\alpha$.
3.1 Overview
CroMe, introduced in this paper, as illustrated in Figure 2, comprises four modules: the encoder, metric learning, Cross-Modal Tri-Transformer Fusion (CMTTF), and fake news detection. The encoder extracts image, text, and integrated encodings. Metric learning enhances intra-modality relationships using the Proxy Anchor method. The CMTTF module fuses these encodings, and the detection module determines whether the content is real or fake.
(a) CMTTF
(b) Inter-Modal Fusion
(c) Tri-Transformer
Figure 4: Overview of the CMTTF architecture, integrating Cross-Modal Fusion and Tri-Transformer modules to process and combine information from text, image, and image-text data. The architecture captures cross-modal correlations and fuses them to enhance feature representation.
3.2 Encoder Module
The encoder layer consists of components for encoding text, images and their combination. These components process input data denoted as $x_{\text{img}}$ for images and $x_{\text{txt}}$ for text.
The image encoder utilizes two parallel methods: Masked Autoencoders (MAE) He et al. (2022) and BLIP2 Image. The pretrained MAE functions as the primary image encoder, producing encodings $Z_{\text{i1}}=f_{\text{MAE}}(x_{\text{img}})$, capturing both global and local features by reconstructing missing data parts. In parallel, BLIP2 Image provides an alternative encoding $Z_{\text{i2}}=f_{\text{BLIP2\_i}}(x_{\text{img}},x_{\text{txt}})$, ensuring that image-specific encodings are derived without textual influence by utilizing a “dummy text” input.
For textual data, BERT and BLIP2 Text encoders are utilized. The pretrained BERT Devlin et al. (2019) generates encodings $Z_{\text{t1}}=f_{\text{BERT}}(x_{\text{txt}})$ by using its bidirectional understanding of text context. Additionally, BLIP2 Text Li et al. (2023) produces encodings $Z_{\text{t2}}=f_{\text{BLIP2\_t}}(x_{\text{img}},x_{\text{txt}})$ by focusing on text data alone, using a zero-filled image tensor to isolate the textual encoding process. The outputs from these encoding processes are concatenated to form integrated image and text encodings, denoted as $Z_{\text{i}}$ and $Z_{\text{t}}$, respectively, for use in subsequent model components.
The BLIP2 model generates combined Image-Text encodings, capturing the cross-modal information between these modalities. This is represented as $Z_{\text{b}}=f_{\text{BLIP2}}(x_{\text{img}},x_{\text{txt}})$, incorporating the integrated features from both image and text inputs.
3.3 Metric Learning
Proxy anchor loss Kim et al. (2020) effectively learns data representations by utilizing distance relationships between data points and proxies. In this work, it captures intra-modal relationships across different modalities, improving the distinction between rumors and non-rumors. The loss function $L(X)$ uses static proxy assignment Movshovitz-Attias et al. (2017), selecting a data point from each class as a proxy. It minimizes the distance between the proxy and same-class data points while maximizing the distance to those from different classes. Modality embeddings $X$ are split into positive ($X^{+}_{p}$) and negative ($X^{-}_{p}$) sets relative to their proxies. The loss is defined as:
$\begin{split}L(X)=&\frac{1}{|P^{+}|}\sum_{p\in P^{+}}\log\bigg{(}1+\sum_{x\in X^{+}_{p}}e^{-\alpha(s(x,p)-\delta)}\bigg{)}\\ &+\frac{1}{|P|}\sum_{p\in P}\log\bigg{(}1+\sum_{x\in X^{-}_{p}}e^{\alpha(s(x,p)+\delta)}\bigg{)}\,,\end{split}$
where $\delta>0$ is the margin parameter (margin) and $\alpha>0$ (alpha) is the scaling parameter.
As shown in Figure 2, the model iteratively trains each modality $*Z_{\text{*}}$ by fixing the parameters of other modalities, allowing it to refine intra-modal relationships without interference. This process is applied to each $Z_{\text{*}}\in\{Z_{\text{i1}},Z_{\text{i2}},Z_{\text{t1}},Z_{\text{t2}},Z_{\text{b}}\}$. Proxy Anchor Loss, illustrated in Figure 3(a), adjusts pull and push forces based on data point proximity, unlike traditional triplet loss Schroff et al. (2015); Peng et al. (2023). Proxy Anchor Loss is computationally efficient as it calculates distances between proxies and data points rather than between individual points, reducing computational load. Applied to text, image, and image-text features, it clusters data points of the same class while separating those of different classes, enhancing fake news detection.
3.4 Cross-Modal Tri-Transformer Fusion
This section introduces Cross-Modal Tri-Transformer Fusion (CMTTF) for integrating and processing image, text, and image-text data. CMTTF combines Cross-Modal Fusion Chen et al. (2022) and Tri-Transformer Choi and Kim (2024), as illustrated in Figure 4(a), with details on each component provided below.
Cross-Modal Fusion
To capture semantic interactions between modalities, the Cross-Modal Fusion module integrates text-image similarities using dot product with softmax normalization and cosine similarity Luo et al. (2018). Given unimodal representations $Z_{\text{i1}}$, $Z_{\text{t1}}$, and additional representations $Z_{\text{i2}}$, $Z_{\text{t2}}$, $Z_{\text{b}}$, the process follows these steps: 1) Inter-modal Fusion and 2) Cosine Similarity.
Inter-modal Fusion integrates semantic interactions between text and image modalities (is illustrated in Figure 4(b)). Correlations between text features $\mathbf{t}$ and image features $\mathbf{i}$ are computed, normalized, and combined into a unified representation $\mathbf{C}_{\text{1}}$, as shown:
$\mathbf{f}_{t\rightarrow i}=\text{Softmax}\left(\frac{\mathbf{t}\cdot\mathbf{i}^{\mathbf{T}}}{\sqrt{d}}\right),$ $\;\mathbf{f}_{i\rightarrow t}=\text{Softmax}\left(\frac{\mathbf{i}\cdot\mathbf{t}^{\mathbf{T}}}{\sqrt{d}}\right),$
$\mathbf{C}_{\text{1}}=\mathbf{f}_{t\rightarrow i}$ $\oplus\mathbf{f}_{i\rightarrow t}\,.$
This inter-modal fusion approach suits models like BERT and MAE, processing unimodal data without embedding normalization. The dot product operation captures direct text-image interactions while preserving scale information, enabling effective multimodal representation.
Cosine Similarity is computed between different combinations of the inputs $Z_{\text{i2}}$, $Z_{\text{t2}}$, and $Z_{\text{b}}$ to measure the relationships between the modalities. The similarities $\mathcal{S}_{\text{ti}}$ ($Z_{\text{t2}}$, $Z_{\text{i2}}$), $\mathcal{S}_{\text{tb}}$ ($Z_{\text{t2}}$, $Z_{\text{b}}$), and $\mathcal{S}_{\text{ib}}$ ($Z_{\text{i2}}$, $Z_{\text{b}}$) are computed using cosine similarity, defined for two vectors $\mathbf{a}$ and $\mathbf{b}$ as:
$\mathcal{S}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\times\|\mathbf{b}\|}\,.$
The combined similarity is computed by weighting individual similarity scores with learnable parameters $w_{\text{i2}}$, $w_{\text{b}}$, $w_{\text{t2}}$, and adding a bias term $b$ to account for modality differences. This similarity is processed through linear layers, ReLU activations, and batch normalization, producing $\mathbf{C}_{\text{2}}$. BLIP2 components use cosine similarity as it effectively captures directional alignment, simplifies the fusion process, and maintains key correlations between modalities Li et al. (2023). The outputs $\mathbf{C}_{\text{1}}$ and $\mathbf{C}_{\text{2}}$ are combined into $Z_{\text{c}}$ then processed through fully connected layers with batch normalization, ReLU activations, and dropout, resulting in the final correlation output used in the subsequent Tri-Transformer.
Tri-Transformer
The text and image embeddings ($Z_{\text{i}}$ and $Z_{\text{t}}$) are passed through their respective projection networks and, along with correlation outputs ($Z_{\text{c}}$), are fed into the Tri-Transformer (Figure 4(c)). This framework is based on the Multi-Modal Tri-Transformer in Choi and Kim (2024). Unlike TT-BLIP’s text-focused cross-modal attention, this method processes text, image, and correlation independently.
$f_{*}=\text{MultiHead}\left(\text{Softmax}\left(\frac{Q_{*}K_{*}^{\mathbf{T}}}{\sqrt{d_{h}}}\right)\times V_{*};\theta_{\text{att}}^{*}\right),$
where $Q_{*},K_{*},V_{*}$ are query, key, and value matrices, and $\theta_{\text{att}}^{*}$ are attention parameters. The outputs processed through Multi-Layer Perceptron (MLP) layers and combined into a unified tensor.
Table 1: Experimental Results for Weibo, Weibo-21 and Politifact Datasets. A dash (‘-’) signifies that results are not provided in their corresponding research paper.
Datasets Methods Accuracy Fake News Real News
Precision Recall F1 Score Precision Recall F1 Score
Weibo EANN Wang et al. (2018) 0.827 0.847 0.812 0.829 0.807 0.843 0.825
MVAE Khattar et al. (2019) 0.824 0.854 0.769 0.809 0.802 0.875 0.837
MPFN Jing et al. (2023) 0.838 0.857 0.894 0.889 0.873 0.863 0.876
Spotfake Singhal et al. (2019) 0.892 0.902 0.964 0.932 0.847 0.656 0.739
SAFE Zhou et al. (2020) 0.762 0.831 0.724 0.774 0.695 0.811 0.748
BDANN Zhang et al. (2020) 0.821 0.790 0.610 0.690 0.830 0.920 0.870
LIIMR Singhal et al. (2022) 0.900 0.882 0.823 0.847 0.908 0.941 0.925
MCAN Wu et al. (2021) 0.899 0.913 0.889 0.901 0.884 0.909 0.897
CAFE Chen et al. (2022) 0.840 0.855 0.830 0.842 0.825 0.851 0.837
FND-CLIP Zhou et al. (2023) 0.907 0.914 0.901 0.908 0.914 0.901 0.907
TT-BLIP Choi and Kim (2024) 0.961 0.979 0.944 0.961 0.944 0.980 0.962
CroMe (Ours) 0.974 0.964 0.984 0.974 0.985 0.966 0.975
Weibo-21 HAN Yang et al. (2016) 0.792 - - 0.792 - - -
Vaibhav(2019) Vaibhav et al. (2019) 0.771 - - 0.771 - - -
FakeFlow Ghanem et al. (2021) 0.781 - - 0.780 - - -
EANN Wang et al. (2018) 0.870 0.902 0.825 0.862 0.841 0.912 0.875
SpotFake Singhal et al. (2019) 0.851 0.953 0.733 0.828 0.786 0.964 0.866
CAFE Chen et al. (2022) 0.882 0.857 0.915 0.885 0.907 0.844 0.876
CroMe (Ours) 0.917 0.944 0.917 0.930 0.880 0.918 0.930
Politifact RoBERTa-MWSS Shu et al. (2004) 0.820 - - 0.820 - - -
SAFE Zhou et al. (2020) 0.874 0.851 0.830 0.840 0.889 0.903 0.896
Spotfake+ Singhal et al. (2020) 0.846 - - - - - -
TM Bhattarai et al. (2022) 0.871 - - - 0.901 - -
LSTM-ATT Lin et al. (2019) 0.832 0.828 0.832 0.830 0.836 0.832 0.829
DistilBert Allein et al. (2021) 0.741 0.875 0.636 0.737 0.647 0.880 0.746
CAFE Chen et al. (2022) 0.864 0.724 0.778 0.750 0.895 0.919 0.907
FND-CLIP Zhou et al. (2023) 0.942 0.897 0.897 0.897 0.960 0.960 0.960
TT-BLIP Choi and Kim (2024) 0.904 0.913 0.724 0.808 0.901 0.973 0.936
CroMe (Ours) 0.933 0.987 0.925 0.955 0.793 0.958 0.868
3.5 Fake News Detector
The fake news detector uses the fusion representation to predict whether news articles are fake or real. This fused representation is processed by a classifier composed of three fully connected layers, ReLU activations, and batch normalization, resulting in a binary classification. The classifier assigns labels of ‘real’ (0) or ‘fake’ (1) to the news content, where $\hat{y}=[\hat{y}_{0},\hat{y}_{1}]$ denotes the probabilities of the news being ‘real’ (0) or ‘fake’ (1). Cross-entropy is used to define the loss function $L(\theta)$, where $\theta$ represents the model parameters.
$L(\theta)=-y\log(\hat{y}_{1})-(1-y)\log(\hat{y}_{0})$
The total loss $L_{\text{total}}$ is defined, where $\alpha$ controls the weight of this metric loss. The optimal value of $\beta$ determined through a grid search over the range [0.1, 1] using a step 0.1. The model aims to minimize the total loss $L_{\text{total}}$ for each news data by learning $\theta$ through back propagation.
$L_{\text{total}}=L(\theta)+\beta\cdot L(X)$
4 Experiments
4.1 Experimental Setup
Datasets
To evaluate the performance of CroMe, three datasets are used: Weibo Jin et al. (2017), Weibo-21 Nan et al. (2021), and Politifact Shu et al. (2020). The Weibo dataset contains 6,137 training articles (2,802 fake and 3,335 real) and 1,685 test articles (833 fake and 852 real). The Weibo-21 dataset includes 4,640 real and 4,487 fake articles, split into training and testing sets with an 8:2 ratio. The Politifact dataset consists of 381 training articles (246 fake and 135 real) and 104 test articles (74 fake and 30 real).
Training Settings
Text encoding utilized the pretrained BERT model Devlin et al. (2019) for Chinese in the Weibo and Weibo-21 datasets and the ”bert-base-uncased” model for the Politifact dataset. Images were resized to $224\times 224$ pixels and encoded using Masked Autoencoders (MAE) He et al. (2022). Image-text pairs were encoded with the pretrained BLIP2 model Li et al. (2023), translating Chinese texts to English via the Google Translation API Johnson (2012). The model employed five Adam optimizers for $Z_{\text{*}}\in\{Z_{\text{i1}},Z_{\text{i2}},Z_{\text{t1}},Z_{\text{t2}},Z_{\text{b}}\}$, with a learning rate of $1\times 10^{-3}$, a batch size of 64, and was trained for 50 epochs. Metric Learning parameters were set as follows: an iterative frequency of 5 epochs per modality, $\alpha=16$, $\delta=0.1$, and $\beta=0.1$, balancing classification and metric losses. These values were determined through preliminary experiments to optimize performance.
4.2 Results and Analysis
CroMe’s performance is evaluated against state-of-the-art models shown in Table 1. The evaluation metrics used include accuracy and precision, recall, and F1 scores for both real and fake news. CroMe achieved the highest accuracy of 0.974 in the experiments using Weibo, surpassing TT-BLIP by 1.3% and FND-CLIP by 6.7%. Similarly, when using Weibo-21, CroMe reached an accuracy of 0.917, outperforming CAFE by 3.5% and EANN by 4.7%. For both datasets, CroMe ranked 1st or 2nd in precision, recall, and F1 scores for both fake and real news. CroMe achieves 0.933 accuracy that closely matches FND-CLIP’s 0.942. The reason may be that the dataset size is too small.
(a) Weibo dataset
(b) Weibo-21 dataset
(c) Politifact dataset
Figure 5: Parameter Analysis for Weibo, Weibo-21 and Politifact dataset using the heatmap.
Table 2: Ablation experimental results of CroMe.
Datasets Modules Accuracy F1 Score
MAE BLIP2Img BERT BLIP2Txt BLIP2Img,Txt CM MT TT Fake News Real News
Weibo 0.961 0.919 0.975
0.950 0.852 0.961
0.949 0.854 0.961
0.971 0.912 0.976
0.971 0.935 0.971
0.959 0.915 0.962
0.906 0.900 0.910
0.890 0.803 0.934
0.974 0.974 0.975
Weibo21 0.903 0.890 0.959
0.882 0.858 0.957
0.890 0.876 0.957
0.897 0.870 0.916
0.899 0.893 0.961
0.899 0.891 0.961
0.895 0.875 0.958
0.888 0.860 0.910
0.917 0.930 0.930
Politifact 0.913 0.927 0.972
0.865 0.891 0.968
0.846 0.839 0.940
0.923 0.933 0.976
0.923 0.942 0.978
0.913 0.915 0.972
0.875 0.790 0.910
0.865 0.877 0.967
0.933 0.955 0.868
(a) w/o image
(b) w/o text
(c) w/o blip2
(d) w/o CMTTF
(e) w/o MT
(f) CroMe
Figure 6: T-SNE visualizations of the features by CroMe and its variants using the Weibo test dataset, with each color representing a distinct label grouping.
CroMe excelled in fake news detection due to three key factors: 1) Advanced multimodal feature extraction through the BLIP2 model enhances CroMe’s capabilities by capturing detailed features from both text and images, allowing for more precise feature extraction. This precision improves the model’s ability to distinguish between real and fake news. 2) The Cross-Modal Tri-Transformer Fusion (CMTTF) computes and integrates cross-modal similarities, enhancing the model’s ability to utilize interactions between different modalities which is crucial for accurate fake news detection. 3) CroMe incorporates a Metric Learning module using the Proxy Anchor method, which focuses on intra-modality relationships. This ensures that features within the same modality are more closely aligned, leading to better representation learning and overall improved performance in detecting fake news. For smaller datasets, future research could explore data augmentation or transfer learning to improve performance.
4.2.1 Proxy Anchor Loss Parameter Analysis
The analysis of the Weibo, Weibo-21, and Politifact datasets under different Proxy Anchor Loss hyperparameters, $\delta$ (margin) and $\alpha$ (alpha), is shown in Figure 5. As the margin increases, performance declines, with larger margins causing greater drops. Performance improves as alpha approaches 16, with the best results at alpha 16. Performance is more sensitive to margin variations than alpha. Alpha was tested at 4, 8, 16, 32, and margin at 0.1, 0.2, 0.3, 0.4. The best performance was achieved with alpha 16 and margin 0.1.
4.3 Ablation
Ablation experiments evaluated the impact of each component in the CroMe model. Table 2 shows the results for the Weibo, Weibo-21, and Politifact datasets. The CroMe variants compared are:
  1. 1.
    Without Image (BLIP2Img + MAE): Removing the image encoders and using only text encoders and other components.
  2. 2.
    Without Text (BLIP2Txt + BERT): Removing the text encoders and using only image encoders and other components.
  3. 3.
    Without BLIP2 (BLIP2Img + BLIP2Txt + BLIP2Img-Txt): Removing all BLIP2 components (text, image and image-text) and using only other encoders.
  4. 4.
    Without BLIP2Img-Txt: Removing the image-text feature component of BLIP2, retaining the individual text and image BLIP2 encoders.
  5. 5.
    Without CM (Cross-Modal): Removing the Cross-Modal Fusion responsible for integrating cross-modal similarities.
  6. 6.
    Without MT (Metric Learning): Removing the Metric Learning module.
  7. 7.
    Without TT (Tri-Transformer): Removing the Tri-Transformer component.
The ablation study highlights three components that most significantly impact accuracy when removed. First, excluding the BLIP2 text components causes a substantial drop in accuracy across all datasets, as they are critical for understanding textual context necessary for identifying fake news. Second, removing the Cross-Modal Fusion (CM) component weakens performance by preventing effective integration of inter-modal fusion and cosine similarity features from text and images. Lastly, the Metric Learning module (MT) is essential for learning distinct features that differentiate fake and real news; its removal reduces accuracy. Overall, uni-modal and cross-modal features, along with CMTTF and Metric Learning, are crucial for the model’s effectiveness in detecting fake news.
4.3.1 T-SNE visualizations
Figure 6 provides a T-SNE Van der Maaten and Hinton (2008) visualization of features before classification, comparing various CroMe settings: CroMe w/o image, CroMe w/o text, CroMe w/o BLIP, CroMe w/o fusion, CroMe w/o metric and the full CroMe model on the Weibo test dataset. Dots of the same color indicate the same label. Figure 6(d) shows that without the Cross-Modal Tri-Transformer Fusion (CMTTF) module, fake and real news instances are not well-separated. Including it improves clustering and distinction between classes. Comparisons of Figures 6(a)6(b)6(c), and 6(f) shows that removing BLIP2 feature extraction results in less clear clustering, highlighting its importance for integrating image, text, and image-text data. Excluding text features also reduces separation in t-SNE plots, showing that image and image-text features alone are less effective. Comparing CroMe w/o MT (Metric Learning, Figure 6(e)) and CroMe (Figure 6(f)), the degree of separation of the sample dots in Figure 6(f) is higher. This indicates that capturing intra-modality relationships ensures effective representation learning within the same modality.
5 Conclusion
This study introduces Cross-Modal Tri-Transformer and Metric Learning (CroMe), a Multimodal Fake News Detection model. CroMe uses Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models (BLIP2) to capture intra-modality and inter-modality relationships. The model includes four main modules: encoders (BERT, BLIP2-text, Masked Autoencoders (MAE), and BLIP2-image), a metric learning module (proxy anchor method), a feature fusion module (Cross-Modal Tri-Transformer Fusion, CMTTF), and a fake news detection module. CroMe uses BLIP2 for semantic information extraction, CMTTF for feature fusion, and the proxy anchor method for metric learning. It improves accuracy by 1.3% on the Weibo dataset and 3.5% on the Weibo-21 dataset compared to previous models. CroMe performed slightly below the state-of-the-art model by 0.9% in the case of Politifact, due to the smaller dataset size.